OpenStudy (anonymous):

How many moles in 360g glucose

OpenStudy (abb0t):

First start with the molecular formula for gluclose: C\(_6\)H\(_{12}\)O\(_6\) Then find the molar mass, which has the units \(\sf \color{}{\frac{g}{mole}}\): \(\sum\)n(mass of each element in C\(_6\)H\(_{12}\)O\(_6\)) the \(n\) the number of carbon, hydrogen and oxygen for EACH atom. So that means (6 \(\times\)mass of 1 carbon atom) + (12 \(\times\) mass of 1 hydrogen atom) +... That is your molar mass. Then similar to how you worked in algebra, cancel out the units. You're given: 360 g of glucose \(\times\) \(\sf \color{}{\frac{mol}{g~of~C_6H_{12}O_6}}\), NOTICE HOW the units cancel out for grams! You're left with only mol. now, do the same thing, to cancel out atoms, by using \(\sf \color{red}{avogadro's~number}\) = \(\sf \color{purple}{6.022 \times 10^{23} \frac{mol}{atom}}\) You're odone.